# Energy-Efficient Split Learning for Fine-Tuning Large Language Models in Edge Networks

### v1

@@ V-A Simulation Setup
We employ a 1B LLaMA 3.2 model with 32-layer transformer decoders [12], and consider 5 mobile devices and one AP with an edge server. These devices within the coverage area of the AP request to access and execute the LLM fine-tuning task at the beginning of each experimental trial. The link bitrate is converted by 5G new radio channel quality indication (CQI) to modulation and coding scheme (MCS) mapping table [13]. The sets of the server and devices are as shown Table I. Here, the computational capabilities of the devices gradually decrease from Device 1 to Device 5. The other main simulation parameters are concluded in Table II.
@@ V-B Simulation Results
In Fig. 3, we first simulate the optimal cut layer and computation resource allocation on these devices. Fig. 3(a) illustrates the optimal cut layer associated with each device, where the optimal cut layer of each device dynamically changes with training rounds due to the dynamic wireless channel. For each device, its optimal cut is either 31 or 1. This is because each transformer layer has the same computation workload and data size as the smashed data, leading to a positive correlation between the cut layer and training delay, and a negative correlation with the server’s energy consumption. In addition, as the computation power of a device decreases, i.e., from Device 1 to 5, its optimal cut layer moves from 31 to 1. Consequently, this necessitates the server to allocate a higher GPU frequency to collectively complete the LLM fine-tuning, as shown in Fig. 3(b). That is because the device with lower computing power prefers to offload the majority of the LLM to the edge server rather than train them locally, which results in reducing the training delay.
@@
Then, we evaluate the fine-tuning performance of the proposed approach in terms of training delay and energy consumption. We compared the proposed approach with two benchmark methods: (i) Server-only, where devices fine-tun the first layer locally, while the server completes the remaining training process. (ii) Device-only, where devices fine-tune the first 31 layers locally, while the server completes the last one-layer training. Following the setting in [14], the channel condition is set to three states, i.e., Good, Normal, and Poor, corresponding to noise spectrum density -166, -163, and -160 dBm/Hz, respectively.
In Fig. 4(a), we shows the training delay during a training round associated with each method, where the proposed scheme gets a performance closer to the server-only method while achieving a 70.8% reduction in training delay compared to the device-only method. Fig. 4(b) depicts the server’s average energy consumption during a training round under each method, illustrating that the proposed approach achieves a 53.1% reduction in the server’s computation consumption compared to the server-only method while consuming a similar computation energy on the server compared to the device-only method.
@@
(b) Average computation energy consumption on the server.
@@ VI Conclusion
We have proposed an energy-efficient SL framework for deploying LLMs in edge networks. We have formulated an optimization problem to achieve the trade-off between the training delay and the server’s energy consumption. To solve this problem, we have designed the CARD algorithm to jointly determine the optimal cut layer and server’s computation resource allocation. Simulation results have demonstrated the effectiveness of the proposed approach. The proposed framework can be applied to accelerate LLM fine-tuning in energy-constrained edge networks. For the future work, we will explore an adaptive strategy to enhance robustness against varying edge network conditions.

### v2

@@ V-A Simulation Setup
We employ a 1B LLaMA 3.2 model with 32-layer transformer decoders [14], and consider 5 mobile devices and one AP with an edge server. These devices within the coverage area of the AP request to access and execute the LLM fine-tuning task at the beginning of each experimental trial. The sets of the server and devices are as shown in Table I. Here, the computational capabilities of the devices gradually decrease from Device 1 to Device 5. The other main simulation parameters are concluded in Table II.
@@ V-B Simulation Results
In Fig. 3, we first simulate these devices’ optimal cut layer and computation resource allocation. Fig. 3(a) illustrates the optimal cut layer associated with each device, where the optimal cut layer of each device dynamically changes with training rounds due to the dynamic wireless channel. For each device, its optimal cut is either 32 or 0. This is because each transformer layer has the same computation workload and data size as the smashed data, leading to a positive correlation between the cut layer and training delay, and a negative correlation with the server’s energy consumption. In addition, as the computation power of a device decreases, i.e., from Device 1 to 5, its optimal cut layer moves from 32 to 0. Consequently, the server allocates a higher GPU frequency to complete the LLM fine-tuning, as shown in Fig. 3(b). That is because the device with lower computing power prefers to offload most of the LLM to the server rather than train them locally, reducing the training delay.
We then evaluate the fine-tuning performance of the proposed approach in terms of training delay and energy consumption, comparing it with two benchmarks: (i) Server-only, where devices fine-tune the embedding module locally, and the server handles the rest. (ii) Device-only, where devices fine-tune the embedding module and transform decoders locally, and the server handles the rest. The channel condition is set to three states, i.e., Good, Normal, and Poor, corresponding to the path-loss exponent 2, 4, and 6, respectively.
In Fig. 4, we present the training delay and the server’s average computational energy consumption during a training round for each method. Compared with the device-only method, the proposed scheme significantly reduces training delay by 70.8%, despite introducing some computational energy consumption on the server. Additionally, while the training delay of the proposed scheme is higher than that of the server-only method, it achieves a remarkable 53.1% reduction in the server’s computational energy consumption.
@@
(b) Average computational energy consumption on the server.
@@ VI Conclusion
We have proposed an energy-efficient SL framework for deploying LLMs in edge networks. We have formulated an optimization problem to achieve the trade-off between the training delay and the server’s energy consumption. To solve this problem, we have designed the CARD algorithm to jointly determine the optimal cut layer and server’s computation resource allocation. Simulation results have demonstrated the effectiveness of the proposed approach. The proposed framework can be applied to accelerate LLM fine-tuning in energy-constrained edge networks. For future work, we will explore an adaptive strategy to enhance robustness against varying edge network conditions.